### v2

# Supplemental Material to ”The Emergence of Vertical Diversity under Disturbance”

Supplemental Material to ”The Emergence of Vertical Diversity under Disturbance”
Yuya Shishikura    Hiroki Ohta
Department of Human Sciences, Obihiro University of Agriculture and Veterinary Medicine, Hokkaido, 080-8555, Japan
(January 12, 2024)
I The equations to determine the stationary solutions of the population dynamics
As mentioned in the main text, based on the model we study, we consider ${\bm{Q}}:=(Q_{1},Q_{2},Q_{3})\in[0,1]^{3}$ where $Q_{h}$ is the density of each state $h$. Then, we may derive the following equation of the population dynamics of ${\bm{Q}}$ for $h=1,3$:
$\frac{dQ_{h}}{dt}=V_{h}({\bm{Q}}),$ (1)
$V_{h}({\bm{Q}})=-\delta_{h,3}rQ_{3}+Q_{2}U\Bigl{(}2\to h\mid{\bm{Q}}\Bigr{)}+\delta_{h,1}r\sum_{h^{\prime}=2}^{3}Q_{h^{\prime}}-Q_{h}U\Bigl{(}h\to 2\mid{\bm{Q}}\Bigr{)}.$
We define $Q_{h}^{*}$ such that $V_{h}({\bm{Q}}^{*})=0$ holds for $h=1,3$.
Let us remind that because of $Q_{1}^{*}+Q_{2}^{*}+Q_{3}^{*}=1$, we have
$Q_{2}^{*}=1-Q_{1}^{*}-Q_{3}^{*},$ (2)
$\sum_{h=1}^{3}hQ_{h}^{*}=-Q_{1}^{*}+Q_{3}^{*}+2.$ (3)
We substitute (2) and (3) into (1) and then obtain the following for $h=1,3$:
$-\delta_{h,3}rQ_{3}^{*}+(1-Q_{1}^{*}-Q_{3}^{*})\frac{(1-r)}{4}\Big{(}1+G_{\beta}^{(h)}(Q_{3}^{*}-Q_{1}^{*})\Big{)}$ (4)
$+\delta_{h,1}r(1-Q_{1}^{*})-Q_{h}^{*}\frac{(1-r)}{4}\Big{(}1-G_{\beta}^{(h)}(Q_{3}^{*}-Q_{1}^{*})\Big{)}=0,$
where
$G_{\beta}^{(h)}(a)=\tanh\beta\Big{(}(\delta_{h,1}-\delta_{h,3})2a+1\Big{)}.$ (5)
I.1 The explicit expressions of $Q_{h}^{*}$ and $\chi_{h}$ at $r\to 0$
Based on (4) for $T>0$, we find the stationary solution of the densities $Q_{1}^{*}$ and $Q_{3}^{*}$ satisfying the following:
$\frac{1}{4}(1-2Q_{1}^{*})(1+\tanh\beta)-\frac{1}{4}Q_{1}^{*}(1-\tanh\beta)=0.$ (6)
Then, taking into account (2), we obtain
$Q_{1}^{*}(T,0)=Q_{3}^{*}(T,0)=\frac{1}{2+\exp(-2\beta)},$ (7)
$Q_{2}^{*}(T,0)=\frac{\exp(-2\beta)}{2+\exp(-2\beta)}.$
Next, we focus on the case of $T=0$ where the following holds:
$\lim_{T\to 0}G_{\beta}^{(h)}(a)=\begin{cases}1&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}>0\\ -1&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}<0\\ 0&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}=0.\end{cases}$ (8)
There are $6$ cases in terms of the possibilities of taking values of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$, each of which can be $-1$, 00, or $1$. Note that if one of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$ is 00, the value of the other is uniquely determined. Indeed, when the two conditions of
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1>0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (9)
hold, based on (4), we obtain
$Q_{3}^{*}=-Q_{1}^{*}+1,$ (10)
where we have used
$\lim_{T\to 0}G_{\beta}^{(h)}(Q_{3}^{*}-Q_{1}^{*})=1,$ (11)
for $h=1,3$. Additionally, the two conditions of (9) lead to
$\frac{1}{4}<Q_{1}^{*}(0,0)<\frac{3}{4},$ (12)
$Q_{3}^{*}(0,0)=1-Q_{1}^{*}(0,0),$
$Q_{2}^{*}(0,0)=0.$
Note that the other $5$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
In order to compute $\chi_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial r}$ for $T>0$, let us perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $r=0$ for $h=1,3$ as follows:
$Q_{h}^{*}(T,r)=Q_{h}^{*}(T,0)+\chi_{h}(T,0)r+O(r^{2}).$ (13)
Then, substituting this form with (7) into (4), for $O(\beta r)\ll 1$ up to $O(r)$ order, we obtain
$-2\chi_{1}-g_{\beta}\chi_{3}+\frac{4g_{\beta}^{\prime}\beta}{2+g_{\beta}}(\chi_{3}-\chi_{1})+\frac{8}{2+g_{\beta}}=0,$ (14)
$-g_{\beta}\chi_{1}-2\chi_{3}+\frac{4g_{\beta}^{\prime}\beta}{2+g_{\beta}}(\chi_{1}-\chi_{3})-\frac{4g_{\beta}}{2+g_{\beta}}=0,$
where
$g_{\beta}=1+\tanh\beta,$ (15)
$g_{\beta}^{\prime}=\frac{1}{\cosh^{2}\beta},$ (16)
with $g_{0}=1\leq g_{\beta}<\lim_{\beta\to\infty}g_{\beta}=2$.
Solving these equations, we finally obtain
$\chi_{1}(T,0)=4\frac{(2+g_{\beta})(4+g_{\beta}^{2})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$ (17)
$\chi_{3}(T,0)=4\frac{-4g_{\beta}(2+g_{\beta})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$
where $g_{\beta}^{\prime}\beta\simeq\beta\exp(-2\beta)\to 0$ as $\beta\to\infty$ and $g_{\beta}^{\prime}\beta\simeq\beta\to 0$ as $\beta\to 0$. Note that $\chi_{2}=-\chi_{1}-\chi_{3}$ by definition.
I.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
Among $6$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, let us consider the situation where the following two conditions hold for $T\ll 1$:
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1=0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0.\end{cases}$ (18)
In this case, we need to take the limit of $T\to 0$ carefully. First, we can perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $T=0$ as
$Q_{h}^{*}(T,r)=Z_{h}+\mu_{h}(0,r)T+O(T^{2}),$ (19)
where $Z_{h}:=Q_{h}^{*}(0,r)$ and $\mu_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial T}$. Substituting this form into (4), we obtain, up to $O(T^{0})$ order,
$\frac{(1-r)}{4}(1-Z_{1}-Z_{3})\Bigl{\{}1+\tanh 2(-\mu_{1}+\mu_{3})\Bigr{\}}$
$-\frac{(1-r)}{4}Z_{1}\Bigl{\{}1-\tanh 2(-\mu_{1}+\mu_{3})\Bigr{\}}+r(1-Z_{1})=0,$ (20)
$\frac{1}{2}(1-r)(1-Z_{1}-Z_{3})-rZ_{3}=0,$ (21)
where the second equation is obtained on the condition of (18) corresponding to
$\lim_{T\to 0}G_{\beta}^{(3)}(Q_{3}^{*}-Q_{1}^{*})\to 1.$ (22)
Combining (21) with the first condition of (18) gives
$Z_{1}=$ $Q_{1}^{*}(0,r)=-\frac{1}{4}r+\frac{3}{4},$ (23)
$Z_{3}=$ $Q_{3}^{*}(0,r)=-\frac{1}{4}r+\frac{1}{4},$
$Q_{2}^{*}(0,r)=\frac{r}{2}.$
Let us move onto (20) into which we substitute (23), leading to
$\tanh 2(-\mu_{1}+\mu_{3})$ $=\frac{-r^{2}-10r+3}{(r+3)(1-r)}=:\theta(r).$ (24)
Since $-1<\tanh 2(-\mu_{1}+\mu_{3})<1$ should hold, $r$ satisfies $0<r<r_{0}=-3+2\sqrt{3}$ due to $\lim_{r\to 0}\theta(r)=1$ and $\lim_{r\to r_{0}}\theta(r)=-1$. Thus, it turns out that (23) as the stationary solution is valid only for $0<r<r_{0}$. Note that $\lim_{r\to 0}Q_{1}^{*}(0,r)=\sup Q_{1}^{*}(0,0)=\frac{3}{4}$, $\lim_{r\to 0}Q_{3}^{*}(0,r)=\inf Q_{3}^{*}(0,0)=\frac{1}{4}$, and $\lim_{r\to 0}Q_{2}^{*}(0,r)=Q_{2}^{*}(0,0)=0$ hold.
Next, among the other $5$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, when the following two conditions
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1<0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (25)
hold, based on (4) we obtain
$-\frac{1}{2}(1-r)Q_{1}^{*}+r(1-Q_{1}^{*})=0,$ (26)
$\frac{1}{2}(1-r)(1-Q_{1}^{*}-Q_{3}^{*})-rQ_{3}^{*}=0,$
where we have used
$\lim_{T\to 0}G_{\beta}^{(1)}(Q_{3}^{*}-Q_{1}^{*})=-1,$ (27)
$\lim_{T\to 0}G_{\beta}^{(3)}(Q_{3}^{*}-Q_{1}^{*})=1.$ (28)
It reads
$Q_{1}^{*}(0,r)=\frac{2r}{1+r},$ (29)
$Q_{3}^{*}(0,r)=\frac{(1-r)^{2}}{(1+r)^{2}},$
$Q_{2}^{*}(0,r)=\frac{2r(1-r)}{(1+r)^{2}}.$
Substituting (29) into the two conditions of (25) gives
$\left\{\begin{array}[]{l}3r^{2}+10r-1>0,\\ r^{2}+6r-3>0,\end{array}\right.$
which lead to $r_{0}<r\leq 1$. Thus, the obtained solution (29) is valid only for $r_{0}<r\leq 1$. Combining two cases of smaller $r$ and larger $r$ than $r_{0}$, the left-sided limit and the right-sided limit of $Q_{h}^{*}$ as $r\to r_{0}$ coincide. Note that the other $4$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
In order to compute the response $\mu_{h}(0,r)$, we can perform Taylor series expansion as
$Q_{h}^{*}(T,r)=Q_{h}^{*}(0,r)+\mu_{h}(0,r)T+O(T^{2}).$ (30)
Substituting this form into (4) with (22) for $0<r<r_{0}$, we obtain, up to $O(T)$ order,
$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})\Bigl{\{}1+\tanh 2(-\mu_{1}+\mu_{3}+O(T))\Bigr{\}}$
$-\frac{(1-r)}{4}Q_{1}^{*}\Bigl{\{}1-\tanh 2(-\mu_{1}+\mu_{3}+O(T))\Bigr{\}}+r(1-Q_{1}^{*})=0,$ (31)
$(1-r)\mu_{1}+(1+r)\mu_{3}=0.$ (32)
Note that the first equation does not give an explicit exact solution up to $O(T)$ order.
Combining (32) with (24), we obtain
$\mu_{1}(0,r)=-\frac{1+r}{4}\text{artanh }\theta(r),$ (33)
$\mu_{3}(0,r)=\frac{1-r}{4}\text{artanh }\theta(r).$
Note that $\mu_{2}=-\mu_{1}-\mu_{3}$ holds by definition, leading to
$\mu_{2}(0,r)=\frac{r}{2}\text{artanh }\theta(r).$ (34)
Thus, it turns out that $|\mu_{h}|\simeq\log(r-r_{0})$ for each $h$.
Let us move onto the case of $r_{0}<r\leq 1$. By using the same Taylor series expansion, instead of (31), we obtain
$-r\mu_{1}-\frac{1}{2}(1-r)\mu_{1}=0.$ (35)
Then, combining this equation with (32) leads to
$\mu_{1}(0,r)=\mu_{3}(0,r)=\mu_{2}(0,r)=0.$ (36)
I.3 The explicit expressions of $Q_{h}^{*}$ and $\nu_{h}$ at $\beta\to 0$
Based on (4), taking into account
$\lim_{\beta\to 0}G_{\beta}^{(h)}(a)=0,$ (37)
we obtain the following:
$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})-\frac{(1-r)}{4}Q_{1}^{*}+r(1-Q_{1}^{*})$ $=0,$ (38)
$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})-\frac{(1-r)}{4}Q_{3}^{*}-rQ_{3}^{*}$ $=0.$
Solving these equations, we obtain
$Q_{1}^{*}(\infty,r)=\frac{5r^{2}+10r+1}{3r^{2}+10r+3},$ (39)
$Q_{3}^{*}(\infty,r)=\frac{(1-r)^{2}}{3r^{2}+10r+3},$
$Q_{2}^{*}(\infty,r)=\frac{-3r^{2}+2r+1}{3r^{2}+10r+3},$
which are shown in Fig. 1(a).
(a)Refer to caption
(b)Refer to caption
Figure 1: (a) $Q_{h}^{*}$ as a function of $r$ for $T=\infty$. (b) $D_{\alpha}$ as a function of $r$ for $T=\infty$.
In order to compute the response $\nu_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial\beta}$, we perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $T=\infty$ as
$Q_{h}^{*}(T,r)=Q_{h}^{*}(\infty,r)+\nu_{h}(\infty,r)\beta+O(\beta^{2}).$ (40)
Substituting this form into (4), we obtain
$-\frac{1-r}{4}(\nu_{1}+\nu_{3})-\frac{1-r}{4}\nu_{1}-r\nu_{1}$ $=0,$ (41)
$-\frac{1-r}{4}(\nu_{1}+\nu_{3})-\frac{1-r}{4}\nu_{3}-r\nu_{3}$ $=0.$
Solving these equations, we obtain
$\nu_{1}=\nu_{3}=0,$ (42)
leading to $\nu_{2}=0$ by definition.
Note that by taking the derivative of the stationary solutions from (39) with respect to $r$, we also have
$\chi_{1}(\infty,0)=\frac{20}{9},$ (43)
$\chi_{3}(\infty,0)=-\frac{16}{9},$ (44)
$\chi_{2}(\infty,0)=-\frac{-4}{9},$ (45)
which are consistent with the expression of (17) in the limit of $\beta\to 0$.
II Hill numbers $D_{\alpha}(r)$ as a function of $r$
Let us define $M_{\alpha}:=\log D_{\alpha}$ and ${\bm{Q}}^{*}:=(Q_{1}^{*},Q_{2}^{*},Q_{3}^{*})\in[0,1]^{3}$. Then, the sign of $\partial_{r}M_{\alpha}$ is the same as that of $\partial_{r}D_{\alpha}$. Indeed, we get the following:
$\partial_{r}M_{\alpha}=\alpha(1-\alpha)^{-1}(\sum_{h}(Q_{h}^{*})^{\alpha})^{-1}F_{\alpha}({\bm{Q}}^{*}),$ (46)
$F_{\alpha}({\bm{Q}}^{*})=\sum_{h}(Q_{h}^{*})^{\alpha-1}\chi_{h}.$ (47)
Thus, the nontrivial part determining the sign of $\partial_{r}D_{\alpha}$ is $F_{\alpha}$, which has key information to find out the local maximum point $r_{\ell}(T,\alpha)$ of $D_{\alpha}$. Note that $F_{\alpha}=0$ at $\alpha=1$ holds due to $\sum_{h}\chi_{h}=0$.
II.1 The existence of a local maximum of $D_{\alpha}$ at $T\to 0$ for any $0<\alpha<\infty$
First, let us consider the case of $r<r_{0}$ and $T\to 0$ where we use equations (23) by putting $r=r_{0}$. Then, as shown in Fig. 2(a), we obtain that $F_{\alpha}<0$ for $\alpha>1$, $F_{\alpha}>0$ for $10.199\leq\alpha<$ 1, and $F_{\alpha}<0$ for $\alpha\leq 0.198$ as $r\to r_{0}$. It means that for $\alpha\geq 0.199$, $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is positive, whereas it is negative for $0<\alpha\leq 0.198$.
Second, let us consider the case of $r>r_{0}$ where we use equations (29) by putting $r=r_{0}$. Then, as shown in Fig. 2(b), it turns out that $F_{\alpha}>0$ for $\alpha>1$ and $F_{\alpha}<0$ for $\alpha<1$. It means that $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is negative for any $\alpha>0$.
Therefore, combining two cases together, we reach the conclusion that there is a local maximum point $r_{\ell}(0,\alpha)>0$ at $T\to 0$ for any $\alpha>0$. The concrete values of $r_{\ell}(0,\alpha)$ are $r_{0}$ for $\alpha\geq 0.199$, and get smaller than $r_{0}$ between $\alpha=0.198$ and $0.199$. Then, $\lim_{\alpha\to 0}r_{\ell}(0,\alpha)$ is numerically estimated as $0.451\cdots$.
(a)Refer to caption
(b)Refer to caption
Figure 2: (a) $\lim_{r\to r_{0}}F_{\alpha}$ as a function of $\alpha$ for $r<r_{0}$. (b) $\lim_{r\to r_{0}}F_{\alpha}$ a function of $\alpha$ for $r>r_{0}$
II.2 The absence of local maxima of $D_{\alpha}$ at $T\to\infty$ for $0<\alpha<\infty$
Suppose the following condition that for a fixed $T$, $\chi_{1}>0$, $\chi_{2}<0$, $\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold for $0<r\leq 1$. On that condition, we can derive that $(Q_{1}^{*})^{\alpha-1}\chi_{1}$ as a term in $F_{\alpha}$ can be larger than $|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ for $\alpha>1$, and smaller than it for $10<\alpha<$ 1. This leads to that for $0<r\leq 1$, $F_{\alpha}>0$ holds for $\alpha>1$, and $F_{\alpha}<0$ holds for $10<\alpha<$ 1, corresponding to that there is no local maximum of $D_{\alpha}$ as a function of $r$ in $0<r\leq 1$ for $\alpha>0$.
The derivation of the above conclusion is as follows. First, it is straightforward to see
$|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ $=(Q_{2}^{*})^{\alpha-1}\chi_{1}+(Q_{3}^{*})^{\alpha-1}|\chi_{3}|-(Q_{2}^{*})^{\alpha-1}|\chi_{3}|$
$=(Q_{3}^{*})^{\alpha-1}\chi_{1}+(Q_{2}^{*})^{\alpha-1}|\chi_{2}|-(Q_{3}^{*})^{\alpha-1}|\chi_{2}|,$ (48)
where we have used $\sum_{h}\chi_{h}=0$. Hence, we obtain
$\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}\leq|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|\leq\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}.$ (49)
It means that for $\alpha>1$, the following holds:
$\chi_{1}(Q_{1}^{*})^{\alpha-1}>\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (50)
which immediately leads to $F_{\alpha}>0$. On the other hand, for $10<\alpha<$ 1, the following holds:
$\chi_{1}(Q_{1}^{*})^{\alpha-1}<\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (51)
which immediately leads to $F_{\alpha}<0$. Thus, we reach the conclusion mentioned above.
For $T=\infty$, (39) directly means that the conditions of $\chi_{1}>0,\chi_{2}<0,\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold. Thus, it leads to the absence of local maxima of $D_{\alpha}$ in $0<r\leq 1$ for any $\alpha>0$ as shown in Fig. 1(b).
Note that the conditions mentioned above can be broken, in particular, for sufficiently small $T$, for which the above results are not applied. Nevertheless, using a nullcline method to compute the stationary solutions of (4), we have numerically estimated that there is $\alpha_{0}(T)>0$ such that for $\alpha>\alpha_{0}(T)$, there are no local maxima of $D_{\alpha}$ for $T\geq 0.26$, whereas for $T\leq 0.25$, there is a local maximum of $D_{\alpha}$ for any $\alpha>0$.
III Monte Carlo sampling
III.1 Updating rules
Let us explain how to obtain sample trajectories by Monte Carlo simulations and compare them with the solutions of the derived population dynamics (1). As Monte Carlo simulations, we perform the following procedures:
  1. Initially, prepare a configuration $\bm{x}$ where the state at each site is a value taken randomly from $\{1,2,3\}$ with uniform distribution. That is, the probability of taking each state is $1/3$.
    Next, repeat the steps from $1$ to $3$:
  2. 1.
    Choose randomly a name $\ell\ (1\leq\ell\leq N)$ of an agent with uniform distribution. That is, the probability of taking each name is $1/N$.
  3. 2.
    With the probability $r$, the chosen agent $\ell$ is disturbed, making the transition from state $x_{\ell}$ to $1$.
  4. 3.
    If the chosen agent $\ell$ is not disturbed, choose either positive or negative direction with probability $1/2$.
    1. (a)
      If positive direction is chosen and $x_{\ell}\neq 3$, make the transition from state $x_{\ell}$ to $x_{\ell}+1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh{\beta\Big{(}2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})}+1\Big{)}\Bigg{)}.$ (52)
    2. (b)
      If negative direction is chosen and $x_{\ell}\neq 1$, make the transition from state $x_{\ell}$ to $x_{\ell}-1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh\beta\Big{(}-2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})+1\Big{)}\Bigg{)}.$ (53)
As shown in Fig. 3, we have confirmed that Monte Carlo simulations have good agreement with the solution of the population dynamics (1).
Figure 3: The Monte Carlo (MC) simulation with $N=1000$ and the solutions of the population dynamics (1). $T=1$, $r=0.2$.
III.2 Finite-size fluctuations
Let us characterize the finite size fluctuations in the density of each state by variance $\sigma_{hh}$, covariance $\sigma_{hh^{\prime}}$, and correlation coefficient $C_{hh^{\prime}}$ for $h,h^{\prime}\in\{1,2,3\}$, which are defined as
$\sigma_{hh^{\prime}}:=\frac{N}{\tau-t_{0}}\sum_{t=t_{0}}^{\tau}\Delta_{h}(t)\Delta_{h^{\prime}}(t),$ (54)
$\Delta_{h}(t):=\hat{Q}_{h}(t)-Q_{h}^{*},$ (55)
$C_{hh^{\prime}}:=\frac{\sigma_{hh^{\prime}}}{\sqrt{\sigma_{hh}}\sqrt{\sigma_{h^{\prime}h^{\prime}}}},$ (56)
where $\hat{Q}_{h}(t)$ is a sample trajectory of density of state $h$ obtained from Monte Carlo simulation at time $t$. The time unit is one Monte Carlo step per site ($N$ steps). We set $t_{0}=100$ because the typical trajectories of $\hat{Q}_{h}(t)$ with $r\geq 10^{-2}$ and $T=10^{-3}$ crosses the stationary solution much before $t=t_{0}=100$. Indeed, for $r\leq 10^{-3}$ and $T=10^{-3}$, the dynamics gets very slow and the typical trajectories do not cross the stationary solutions by $t=t_{0}=100$.
Figure 4: The variance $\sigma_{hh}$ and correlation coefficients $C_{hh^{\prime}}$ averaged over $10$ samples with different trajectories. $T=10^{-3}$, $t_{0}=100$, $\tau=150$, $N=10^{2},10^{3},10^{4}$.
As shown in Fig. 4, we have performed Monte Carlo simulations with different systems size $N$. If we straightforwardly extrapolate the behaviors for larger system size, there seem to be singular behaviors in the three types of variance and the three types of correlation function in the large size limit near $r=r_{0}$ for $r<r_{0}$. In particular, it is notable to point out that three correlation coefficients in the large size limit seem to have perfect correlation, corresponding to $|C_{hh^{\prime}}|=1$, for only $r<r_{0}$.
Thus, the results obtained by Monte Carlo simulations suggest that there is a universal curve presumably independent of $N$. Further, in the larger system sizes, it is more obvious that $C_{hh^{\prime}}$ for $r<r_{0}$ show U-shaped curves. It is notable that the bottoms of the U-shaped curves are close to $r=-5+2\sqrt{7}\simeq 0.291\cdots$, which is the point where $\mu_{h}$ changes its sign. Nevertheless, more works with sufficiently large samples need to be done to obtain more convincing data, which remain to be found as another work in the future.

### v4

@@
-Supplemental Material to ”The Emergence of Vertical Diversity under Disturbance”
-Supplemental Material to ”The Emergence of Vertical Diversity under Disturbance”
-Yuya Shishikura    Hiroki Ohta
-Department of Human Sciences, Obihiro University of Agriculture and Veterinary Medicine, Hokkaido, 080-8555, Japan
-(January 12, 2024)
-I The equations to determine the stationary solutions of the population dynamics
+Emergence of vertical diversity under disturbance
+Abstract
+We propose a statistical physics model of a neutral community, where each agent can represent identical plant species growing in the vertical direction with sunlight in the form of rich-get-richer competition.
+Disturbance added to this ecosystem, which makes an agent restart from the lowest growth level, is realized as a stochastic resetting.
+We show that in this model for sufficiently strong competition, vertical diversity characterized by a family of Hill numbers robustly emerges as a local maximum at intermediate disturbance.
+I Introduction
+Organisms often compete with other organisms and those could coexist as a whole ecosystem.
+The theoretical frameworks to understand the mechanism of such coexistence in ecology have been developed based on Lotka-Volterra equation HS .
+The developed theories have continued to overcome the obstacles to clarify unexplored coexistence mechanisms Tilman82 ; Chesson00 ; Kessler15 ; Obuchi16 ; Bunin17 ; Monasson17 ; Biroli18 ; Cui20 ; Gupta21 ; Barbier21 ; Ros23 .
+One of the important topics in coexistence theory is diversity Magurran10 , which is characterized by various indexes such as Hill numbers equivalent to Rényi entropy Leinster21 .
+Diversity of an ecosystem is largely affected by so-called disturbance Viljur22 .
+Disturbance is a general term referring to the effects led by the environment outside of an ecosystem, which causes a certain loss of mass of organisms in the ecosystem.
+There is a qualitative hypothesis that high diversity is achieved at intermediate disturbance, which is called the Intermediate Disturbance Hypothesis (IDH) Grime73 ; Connell78 .
+Various mathematical models have been found to show what IDH implies Petraitis89 ; Caswell91 ; Kondoh01 ; Roxburgh04 ; Svensson12 ; Hunt21 and also a proposal of updating such a hypothesis has been posed Fox13 .
+The concept of diversity in ecology is not only used about species, but also used about the spatial structure of an ecosystem such as the height of plants Franklin2002 .
+This structural diversity has been discussed in strong relation to species diversity Inoue01 ; Ehbrecht21 .
+Indeed, in German forests, structural diversity in terms of canopy height, which is so-called vertical diversity, has been discovered to get the highest at an intermediate disturbance Senf20 .
+Mathematical models have been also developed, which focus on height statistics of plants Hara84 ; Yokozawa99 ; Zavala07 ; Strigul08 ; Purves08 ; Kohyama09 ; Cammarano11 ; Kohyama12 .
+Nevertheless, IDH in terms of vertical diversity has hardly been mentioned from the perspective of such theoretical models.
+In this paper, we develop a theoretical framework of statistical physics models, where one can quantitatively investigate the relationship among competition, disturbance, and vertical diversity in a community of identical species.
+The approach taken hereafter can also be regarded as a proposal on a unique branch of neutral theories in community ecology MacArthur ; Caswell76 ; Hubbell ; VH2003 ; MAS2004 ; Beeravolu09 ; Volkov03 .
+Indeed, it has already been discussed how theoretical development can go beyond neutral theories Azaele16 and how the neutral theories ignoring spatially heterogeneous effects can be extended to include such spatial effects Pigolotti18 .
+II Model
+In this paper, we consider competition among plants for sunlight, which often takes the form of competition such that taller plants are more likely to occupy sunlight and then to get taller, which is the so-called rich-get-richer form of competition Merton68 .
+As preliminaries, suppose that for a given height distribution of plants, we classify it into three height levels by putting two thresholds of the average plus or minus standard deviation.
+Specifically, let $N\in\mathbb{Z}$ be the total number of agents, $i\in\{1,\dots,N\}$ be the name of each agent, and $x_{i}\in\{1,2,3\}$ be the state as the relative height of agent $i$.
+We use a notation $\bm{x}:=(x_{1},x_{2},\cdots,x_{N})\in\{1,2,3\}^{N}$ as a state of the whole system.
+Based on the above considerations, we introduce a stochastic process where the transition rate representing rich-get-richer competition is induced by a simple energy function $E(\bm{x})$.
+For simplicity, the energy function is required to have the following two conditions: (i) The energy is invariant in terms of replacing the state of maximum $3$ by the state of minimum $1$ in the whole system: $E(\bm{x})=E(\bm{x}^{\prime})$ for $x_{i}^{\prime}=-(x_{i}-2)+2$.
+(ii) In the configuration achieving the minimum energy, the number of state $2$ is zero and also the number of state $3$ minus the number of state $1$ is either zero or ±$\pm 1$.
+Under these two conditions, we consider the following energy function $E(\bm{x})$ in which each agent interacts with agents at neighboring sites determined by the edges of a graph $\mathcal{G}$,
+$E(\bm{x}):=-\frac{1}{N-1}\sum_{i\in V(\mathcal{G})}\sum_{j\in B_{i}(\mathcal{G})}(x_{i}-x_{j})^{2},$ (1)
+where $V(\mathcal{G})$ is the set of all sites in graph $\mathcal{G}$ and $B_{i}(\mathcal{G})$ for $1\leq i\leq N$ is the set of all the nearest neighbor sites of site $i$.
+Indeed, this energy function (1) satisfies (i).
+Further, (1) satisfies (ii) because the minimum value of the energy $-2N$ at the leading order of $N$ is realized when $x_{i}=1$ or $3$ for any $i$ on the condition that the number of state $1$ minus the number of state $3$ is either 00 for even $N$ or ±$\pm 1$ for odd $N$.
+For simplicity, we focus on only the case that $\mathcal{G}$ is the complete graph having edges by which any pair of two sites is directly connected.
+Next, let us define $f_{i}^{\pm}$ such that $f^{\pm}_{i}\bm{x}:=(x_{j}\pm\delta_{i,j})^{N}_{j=1}$ for only the case that $1\leq f_{i}^{\pm}x_{i}\leq 3$, where $\delta_{i,j}$ is the Kronecker delta with $\delta_{a,b}=1$ for $a=b$, otherwise 00.
+Then, we introduce a Markov process in continuous time $t$ as follows.
+Let us consider a transition rate $T_{0}(\bm{x}\to f^{\pm}_{i}\bm{x})$ from $\bm{x}$ to $f^{\pm}_{i}\bm{x}$ is
+$T_{0}(\bm{x}\to f^{\pm}_{i}\bm{x})=\frac{1}{4}\biggl{[}1+\tanh\bigg{(}-\frac{1}{2}\beta\Delta E(\bm{x}\to f^{\pm}_{i}\bm{x})\bigg{)}\biggr{]},$ (2)
+where $0\leq\beta$ and $\Delta{E(\bm{x}\to f^{\pm}_{i}\bm{x})}:=E(f^{\pm}_{i}\bm{x})-E(\bm{x})$
+$=2\biggl{\{}\mp 2\biggl{(}x_{i}-\frac{1}{N-1}\sum_{j\neq{i}}x_{j}\biggr{)}-1\biggr{\}}.$ (3)
+Figure 1: Schematic diagram for the transition rules between two states for the case of three plants $i,j,k$.
+Under sunlight competition, the transitions from $x_{j}=2$ to $3$ or $1$ and the reverse transitions as described by dotted blue arrows may occur.
+Due to disturbance, the transitions from $x_{k}=3$ and $x_{j}=2$ to $1$ described as filled red arrows may occur.
+II.1 Stochastic resetting as disturbance
+Next, as a disturbance added to this system, we consider a transition from the current state of an agent to its lowest growth state.
+This transition can be regarded as one class of stochastic resetting Shamik23 .
+Concretely, let us consider stochastic resetting from state $x_{i}\geq 2$ to state $1$ with a disturbance rate $R(\bm{x}\to d_{i}\bm{x})$ where $d_{i}\bm{x}:=(x_{j}(1-\delta_{i,j})+\delta_{i,j})^{N}_{j=1}$ in the following manner:
+$R(\bm{x}\to d_{i}\bm{x})=r,$ (4)
+where $0\leq r\leq 1$.
+Combining this stochastic resetting with the above Markov process representing competition, the transition rate $T(\bm{x}\to\bm{x^{\prime}})$ of the stochastic process that we consider, as schematically illustrated in Fig.
+1, is totally
+$T(\bm{x}\to\bm{x^{\prime}})=$ $(1-r)\sum_{1\leq i\leq N}\sum_{s=\pm}\delta_{\bm{x^{\prime}},f^{s}_{i}\bm{x}}T_{0}(\bm{x}\to f^{s}_{i}\bm{x})$
+$+\sum_{1\leq i\leq N}\delta_{\bm{x^{\prime}},d_{i}\bm{x}}R(\bm{x}\to d_{i}\bm{x}).$ (5)
+Note that the dynamics governed by the introduced rich-get-richer form of competition and disturbance is neutral in the following sense: the system has essentially the same dynamics governed by the rules of competition and disturbance even if any pairs of individuals are exchanged.
+Next, let $P_{t}(\bm{x})$ be the probability that the state is $\bm{x}$ at time $t$. Then, transition rate (II.1) leads to the following master equation:
+$\frac{dP_{t}(\bm{x})}{dt}=\sum_{\bm{x^{\prime}}\neq\bm{x}}\Bigg{(}P_{t}(\bm{x^{\prime}})T(\bm{x^{\prime}}\to\bm{x})-P_{t}(\bm{x})T(\bm{x}\to\bm{x^{\prime}})\Bigg{)}.$ (6)
+The stationary distribution $P_{\rm st}(\bm{x})$ is defined such that $P_{t}(\bm{x})=P_{\rm st}(\bm{x})$ leads to $\frac{dP_{t}(\bm{x})}{dt}=0$.
+Indeed, the canonical distribution $P_{\rm can}(\bm{x}):=\frac{1}{Z_{N}(\beta)}\exp(-\beta E(\bm{x}))$, where $Z_{N}(\beta)=\sum_{\bm{x}}\exp(-\beta E(\bm{x}))$, satisfies the following detailed balanced condition
+$P_{\rm can}(\bm{x})T_{0}(\bm{x}\to f^{\pm}_{i}\bm{x})=P_{\rm can}(f^{\pm}_{i}\bm{x})T_{0}(f^{\pm}_{i}\bm{x}\to\bm{x}).$ (7)
+Hence, the canonical distribution is the stationary distribution satisfying Eq. (6) with $r=0$.
+If the disturbance rate holds $r>0$, this stochastic process involves irreversible processes caused by disturbance, because the transition from state $1$ to state $3$ does not exist but the reverse exists.
+Thus, the detailed balanced condition can not be satisfied under the presence of disturbance, meaning that $P_{\rm st}(\bm{x})$ no longer follows the canonical distribution determined by the energy function (1).
+III On population dynamics
+Figure 2: (a) The density $Q_{h}^{*}(T,0)$ as a function of $T$.
+At $T=0$, $Q_{1}^{*}$ and $Q_{3}^{*}$ have an infinite number of solutions satisfying $41/4<Q_{1}^{*},Q_{3}^{*}<3/$ 4 and $Q_{2}^{*}=0$.
+At $T\to\infty$, $Q_{h}^{*}(T,0)\to 1/3$ for any $h$.
+(b) $r$-susceptability $\chi_{h}(T,0)$ at $r=0$ as a function of $T$.
+$\chi_{h}(T,0)$ shows an exponential divergence as $T\to 0$.
+III.1 Derivation of population dynamics
+Although the master equation has the complete information about the stationary distribution, in order to avoid dealing with $3^{N}$ states, we consider the time evolution of the density of each of three different states in the limit of large $N$.
+We show how to derive such a set of closed equations in the following.
+First, let $Q_{h}$ be the density of states $h\in\{1,2,3\}$, and be ${\bm{Q}}:=(Q_{1},Q_{2},Q_{3})\in[0,1]^{3}$.
+Then we focus on the transition rate $U\Bigl{(}h\to h\pm 1\mid{\bm{Q}}\Bigr{)}$ from state $h$ to $h\pm 1\in\{1,2,3\}$ of an agent for given ${\bm{Q}}$ in the limit of large $N$.
+Taking into account (2) and (3), $U$ turns out to be equal to
+$\frac{(1-r)}{4}\biggl{[}1+\tanh\bigg{(}-\beta\Delta e_{\pm}\Big{(}h\mid{\bm{Q}}\Big{)}\bigg{)}\biggr{]},$ (8)
+where
+$\Delta e_{\pm}\Bigl{(}h\mid{\bm{Q}}\Bigr{)}=\biggl{\{}\mp 2\biggl{(}h-\sum_{h^{\prime}=1}^{3}h^{\prime}Q_{h^{\prime}}\biggr{)}-1\biggr{\}}.$ (9)
+Note that it is assumed that the replacement of $\sum_{i}x_{i}/N$ by $\sum_{h^{\prime}=1}^{3}h^{\prime}Q_{h^{\prime}}$ in the limit of $N\to\infty$ holds, which can be regarded as self-averaging.
+Taking into account disturbance rate (4), the time evolution of each density $Q_{h}$ for $h=1$ and $3$ is exactly described as $\frac{dQ_{h}}{dt}=V_{h}({\bm{Q}})$, which is explicitly:
+$\frac{dQ_{h}}{dt}$ $=-\delta_{h,3}rQ_{3}+Q_{2}U\Bigl{(}2\to h\mid{\bm{Q}}\Bigr{)}$ (10)
+$+\delta_{h,1}r\sum_{h^{\prime}=2}^{3}Q_{h^{\prime}}-Q_{h}U\Bigl{(}h\to 2\mid{\bm{Q}}\Bigr{)}.$
+III.2 Densities and susceptibility in stationary states
+Before discussing vertical diversity, we try to capture the typical behaviors of the model in the parameter space of disturbance characterized by $r$ and interaction strength characterized by $T=\beta^{-1}$.
+Let us compute the stationary solutions of the derived population dynamics (10).
+The density of each state $h$ in a stationary solution satisfying $V_{h}({\bm{Q}}^{*})=0$ is described by $Q_{h}^{*}(T,r)$ as a function of two parameters of $T=\beta^{-1}$ and $r$.
+Note that because of $Q_{1}^{*}+Q_{2}^{*}+Q_{3}^{*}=1$, we have $Q_{2}^{*}=1-Q_{1}^{*}-Q_{3}^{*}$ and also $\sum_{h=1}^{3}hQ_{h}^{*}=-Q_{1}^{*}+Q_{3}^{*}+2$, by which Eq.
+(10) is written as two coupled equations depending on only $Q_{1}^{*}$ and $Q_{3}^{*}$.
+Let us consider the case of $r=0$ for any $T$. As shown in Fig. 2(a), it turns out that there is a symmetric solution
+$Q_{1}^{*}(T,0)=Q_{3}^{*}(T,0)=(2+\exp(-2\beta))^{-1},$ (11)
+leading to $Q_{h}^{*}(T,0)\to 1/3$ for any $h$ at $T\to\infty$. At $T=0$, in addition to this symmetric solution, we obtain the other asymmetric solutions as follows:
+$\frac{1}{4}<Q_{1}^{*}(0,0)=1-Q_{3}^{*}(0,0)<\frac{3}{4},$ (12)
+which means that infinite number of the solutions $Q_{h}^{*}$ exist under condition (12).
+Figure 3: (a) The density $Q_{h}^{*}(0,r)$ as a function of disturbance rate $r$.
+$Q_{h}^{*}$ for any $h$ are not differentiable at $r=r_{0}$.
+(b) $T$-susceptibility $\mu_{h}(0,r)$ as a function of disturbance rate $r$.
+$\mu_{h}(0,r)$ shows logarithmic divergences as $r\to 0$ for only $h=1,3$ and $r\to r_{0}$ for any $h$.
+Next, let us consider a response of the density against $r$ around $r\ll 1$, which is characterized by $r$-susceptibility
+$\chi_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial r}$ (13)
+for any $h$. As shown in Fig. 2(b), using Taylor series expansion, we straightforwardly obtain
+$\chi_{1}(T,0)=4\frac{(2+g_{\beta})(4+g_{\beta}^{2})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$ (14)
+$\chi_{3}(T,0)=4\frac{-4g_{\beta}(2+g_{\beta})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$ (15)
+where $g_{\beta}=1+\tanh\beta$ with $g_{0}=1\leq g_{\beta}<2=\lim_{\beta\to\infty}g_{\beta}$ and $g^{\prime}_{\beta}=(\cosh\beta)^{-2}$.
+Note that $\chi_{2}=-\chi_{1}-\chi_{3}$ by definition.
+It is notable that $\chi_{h}$ for each $h$ shows an exponential divergence as $\beta\to\infty$ ($T\to 0$).
+Let us move on to the case of $T\ll 1$ for any $r>0$. Similar to the case with $r=0$, we consider a response of the density against $T$ around $T\ll 1$, which is characterized by $T$-susceptibility
+$\mu_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial T}$ (16)
+for any $h$. As shown in Fig. 3, $Q_{h}^{*}$ and $\mu_{h}$ in the limit of $T\to 0$ are computed in the following. First, for $r_{0}=-3+2\sqrt{3}<r\leq 1$, we obtain
+$Q_{1}^{*}(0,r)$ $=\frac{2r}{1+r},$ (17)
+$Q_{3}^{*}(0,r)$ $=\frac{(1-r)^{2}}{(1+r)^{2}},$ (18)
+and $\mu_{h}(0,r)=0$ for any $h$.
+[…]
+$Q_{1}^{*}(0,r)=Q_{3}^{*}(0,r)+\frac{1}{2}=-\frac{1}{4}r+\frac{3}{4},$ (19)
+$\mu_{1}(0,r)=-\frac{1+r}{1-r}\mu_{3}(0,r)=-\frac{1+r}{4}\text{artanh }\theta(r),$ (20)
+where
+$\theta(r)=\frac{-r^{2}-10r+3}{(1-r)(3+r)}.$ (21)
+Note that $\lim_{r\to r_{0}}\theta(r)=-1<\theta(r)<1=\lim_{r\to 0}\theta(r)$ for $0<r<r_{0}$, and $\mu_{2}=-\mu_{1}-\mu_{3}$ by definition.
+It is notable that the average height defined by $\sum_{h=1}^{3}hQ_{h}^{*}$ is equal to $3/2$ independent of $r$ for $0<r<r_{0}$ with $T\to 0$.
+$\mu_{h}$ for $h=1,3$ as $r\to 0$ and for any $h$ as $r\to r_{0}$ show logarithmic divergences.
+Further, finite size fluctuations have been found to show distinct behaviors between lower $r$ and higher $r$ compared to $r_{0}$ (see Appendix.
+B).
+Figure 4: $Q_{h}^{*}$ as a function of $r$ for $\beta=0$. Note that $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ and $\chi_{1}>0,\chi_{2}<0,\chi_{3}<0$ hold for $0<r$.
+For $T\to\infty$, as shown in Fig. 4, we can straightforwardly obtain the analytical form of $Q_{h}^{*}(\infty,r)$ and $\chi_{h}(\infty,r)$ as
+$Q_{1}^{*}(\infty,r)=\frac{5r^{2}+10r+1}{3r^{2}+10r+3},$ (22)
+$Q_{3}^{*}(\infty,r)=\frac{(1-r)^{2}}{3r^{2}+10r+3},$ (23)
+$Q_{2}^{*}(\infty,r)=\frac{-3r^{2}+2r+1}{3r^{2}+10r+3},$ (24)
+which satisfy $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ and $\chi_{1}>0,\chi_{2}<0,\chi_{3}<0$ for any $r>0$ as shown in Fig.
+4.
+Note that $\nu_{h}(\infty,r):=\frac{\partial Q_{h}^{*}}{\partial\beta}=0$ holds for any $h$.
+IV Diversity indexes
+Let us focus on Hill numbers for $\alpha>0$ and $\alpha\neq 1$,
+$D_{\alpha}:=\biggl{(}\sum_{h=1}^{3}Q_{h}^{\alpha}\biggl{)}^{1/(1-\alpha)},$ (25)
+which satisfies the replication principle Leinster21 and quantifies the vertical diversity of the model with $1\leq D_{\alpha}\leq 3$. Further, we define
+$D_{1}:=\lim_{\alpha\to 1}D_{\alpha}=\exp(-\sum^{3}_{h=1}Q_{h}^{*}\log_{e}{Q_{h}^{*}}),$ (26)
+which is equivalent to the exponential of Shannon entropy, where $e$ is Napier’s constant.
+In Fig.
+5, $D_{\alpha}$ for $\alpha=0.1,1,10,100$ are shown at $T\to 0$ and $T\to\infty$.
+Whereas the diversity indexes at $T\to\infty$ show monotonically decreasing behaviors, those at $T\to 0$ robustly get a maximum as a function of a single variable $r$ at an intermediate disturbance $r=r_{0}$ for $\alpha\geq 0.199$.
+In order to elaborate on Hill numbers, we discuss theoretical results in the cases of $T\to\infty$ and $T\to 0$ in the following subsection A.
+Then, in the following subsection B, we present numerical computation of Hill numbers for more general values of $T$.
+(b) Refer to caption
+Figure 5: Diversity indexes $D_{\alpha}(r)$ as a function of $r$ for $T\to\infty$ (a) and $T\to 0$.
+(b) $\alpha=0.1,1,10,100$.
+$D_{\alpha}(r)$ do not have local maxima as a function of $r$ at $T\to\infty$.
+On the contrary, at $T\to 0$, $D_{\alpha}(r)$ have local maxima at intermediate values of disturbance $r$ for any $\alpha$.
+IV.1 Hill numbers $D_{\alpha}(r)$ as a function of $r$
+In order to discuss how the diversity indexes behave for general values of $\alpha$, let us define $M_{\alpha}:=\log D_{\alpha}$ and ${\bm{Q}}^{*}:=(Q_{1}^{*},Q_{2}^{*},Q_{3}^{*})\in[0,1]^{3}$.
+Then, the sign of $\partial_{r}M_{\alpha}$ is the same as that of $\partial_{r}D_{\alpha}$.
+Indeed, we get the following:
+$\partial_{r}M_{\alpha}=\alpha(1-\alpha)^{-1}(\sum_{h}(Q_{h}^{*})^{\alpha})^{-1}F_{\alpha}({\bm{Q}}^{*}),$ (27)
+$F_{\alpha}({\bm{Q}}^{*})=\sum_{h}(Q_{h}^{*})^{\alpha-1}\chi_{h}.$ (28)
+Thus, the nontrivial part determining the sign of $\partial_{r}D_{\alpha}$ is $F_{\alpha}$, which has key information to find out the local maximum point $r_{\ell}(T,\alpha)$ of $D_{\alpha}$.
+Note that $F_{\alpha}=0$ at $\alpha=1$ holds due to $\sum_{h}\chi_{h}=0$.
+IV.1.1 The absence of local maxima of $D_{\alpha}$ at $T\to\infty$ for $0<\alpha<\infty$
+Suppose the following condition that for a fixed $T$, $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ and $\chi_{1}>0$, $\chi_{2}<0$, $\chi_{3}<0$ hold for $0<r\leq 1$.
+On that condition, we can derive that $(Q_{1}^{*})^{\alpha-1}\chi_{1}$ as a term in $F_{\alpha}$ can be larger than $|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ for $\alpha>1$, and smaller than it for $10<\alpha<$ 1.
+This leads to that for $0<r\leq 1$, $F_{\alpha}>0$ holds for $\alpha>1$, and $F_{\alpha}<0$ holds for $10<\alpha<$ 1, corresponding to that there is no local maximum of $D_{\alpha}$ as a function of $r$ in $0<r\leq 1$ for $\alpha>0$.
+Figure 6: (a) $\lim_{r\to r_{0}}F_{\alpha}$ as a function of $\alpha$ for $r<r_{0}$ and (b) for $r>r_{0}$.
+Reminding of (27), it turns out that $\partial_{r}\log D_{\alpha}$ are negative for any $\alpha$ in the case of $r>r_{0}$ and for, at least, $\alpha\geq 0.199$ in the case of $r<r_{0}$.
+On the contrary, $\partial_{r}\log D_{\alpha}$ are positive for, at least, $\alpha\leq 0.198$ in the case of $r<r_{0}$, which means that the local maxima for $\alpha\leq 0.198$ are not located at $r=r_{0}$.
+The derivation of the above conclusion is as follows. First, it is straightforward to see
+$|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$
+$=(Q_{2}^{*})^{\alpha-1}\chi_{1}+(Q_{3}^{*})^{\alpha-1}|\chi_{3}|-(Q_{2}^{*})^{\alpha-1}|\chi_{3}|$
+$=(Q_{3}^{*})^{\alpha-1}\chi_{1}+(Q_{2}^{*})^{\alpha-1}|\chi_{2}|-(Q_{3}^{*})^{\alpha-1}|\chi_{2}|,$ (29)
+where we have used $\sum_{h}\chi_{h}=0$. Hence, we obtain
+$\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}$
+$\leq|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$
+$\leq\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}.$ (30)
+It means that for $\alpha>1$, the following holds:
+$\chi_{1}(Q_{1}^{*})^{\alpha-1}>\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (31)
+which immediately leads to $F_{\alpha}>0$. On the other hand, for $10<\alpha<$ 1, the following holds:
+$\chi_{1}(Q_{1}^{*})^{\alpha-1}<\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (32)
+which immediately leads to $F_{\alpha}<0$. Thus, we reach the conclusion mentioned above.
+Note that for $T=\infty$, (22),(23), and (24) directly mean that the conditions of $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ and $\chi_{1}>0,\chi_{2}<0,\chi_{3}<0$ hold.
+Thus, it leads to the absence of local maxima of $D_{\alpha}$ in $0<r\leq 1$ for any $\alpha>0$ as shown in Fig.
+5(a).
+IV.1.2 The existence of a local maximum of $D_{\alpha}$ at $T\to 0$ for $0<\alpha<\infty$
+First, let us consider the case of $r<r_{0}$ and $T\to 0$ where we use Eq.
+(19) by putting $r=r_{0}$.
+Then, as shown in Fig.
+6(a), we obtain that $F_{\alpha}<0$ for $\alpha>1$, $F_{\alpha}>0$ for $10.199\leq\alpha<$ 1, and $F_{\alpha}<0$ for $\alpha\leq 0.198$ as $r\to r_{0}$.
+It means that for $\alpha\geq 0.199$, $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is positive, whereas it is negative for $0<\alpha\leq 0.198$.
+Second, let us consider the case of $r>r_{0}$ where we use Eqs.
+(17) and (18) by putting $r=r_{0}$.
+Then, as shown in Fig.
+6(b), it turns out that $F_{\alpha}>0$ for $\alpha>1$ and $F_{\alpha}<0$ for $\alpha<1$.
+It means that $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is negative for any $\alpha>0$.
+Therefore, combining two cases together, we reach the conclusion that there is a local maximum point $r_{\ell}(0,\alpha)>0$ at $T\to 0$ for any $\alpha>0$.
+The concrete values of $r_{\ell}(0,\alpha)$ are $r_{0}$ for $\alpha\geq 0.199$, and get smaller than $r_{0}$ between $\alpha=0.198$ and $0.199$.
+Then, $\lim_{\alpha\to 0}r_{\ell}(0,\alpha)$ is numerically estimated as $0.451\cdots$.
+Figure 7: The local maximum points $r_{\ell}(T,\alpha)$ of $D_{\alpha}(r)$ depending on $(T,\alpha)$ for $Q_{h}^{*}$.
+$r_{\ell}(0,0.1)=0.457\cdots<r_{0}=0.464\cdots$.
+The data points of $r_{\ell}(T,\alpha)$ are estimated within the numerical resolution by a nullcline method where the increment of $r$ and $T$ is $10^{-2}$.
+$T=0.25$ is around the value, above which local maxima for certain values of $\alpha$ do not appear.
+IV.2 Computation of Hill numbers by nullcline method
+Next, we have performed the nullcline method for computing ${\bm{Q}}^{*}$ to estimate the local maximum of $D_{\alpha}(r)$ as a function of $r$ for general values of $T$ and $\alpha$.
+As shown on the $(T,r)$-parameter plane in Fig.
+7, whereas $r=0$ could also get the local or global maximum point for any $T>0$, local maximum points $r_{\ell}(T,\alpha)>0$ of $D_{\alpha}$ as a function of single variable $r$ are determined by $\partial_{r}D_{\alpha}(r_{\ell}(T,\alpha))=0$.
+Indeed, there is a point around $T=0.25$ below which there exists a local maximum point for any $\alpha>0$.
+On the contrary, for $T\geq 0.26$, $r_{\ell}(T,\alpha)$ does not exist for all the values of $\alpha$ higher than a sufficiently large $\alpha_{0}(T)$.
+Related to this phenomenon, we have also found that the local minima of $Q_{1}^{*}$ as a function of single variable $r$, which are observed for $T\leq 0.25$, disappear for $T\geq 0.26$.
+Further, the local maxima of $Q_{2}$ exist for $T\leq 2.58$, and it seems to disappear between $T=2.58$ and $T=2.60$, which is very close to the lower limit of the absence of local maxima of $D_{\alpha}$ for any $\alpha$.
+V Concluding remarks
+In this paper, we have proposed a neutral community model of identical plant species with rich-get-richer competition, which robustly exhibit emergent vertical diversity that has a local maximum at an intermediate disturbance for sufficient strong competition.
+This property observed in the proposed model is qualitatively consistent with the property observed in German forests Senf20 .
+It remains for future studies to perform more detailed comparisons with the properties observed in real ecosystems.
+It is intriguing to consider how to blend the effects of migration and death process from classical neutral models Hubbell ; VH2003 ; MAS2004 with the model in this paper.
+Future works in that direction could provide unique insights into neutral species abundance distributions into vertically-structured local communities.
+The proposed neutral model can be easily generalized into a non-neutral model with multiple species by introducing species dependence into transition rates representing competition or disturbance.
+In another direction, one can study spatial effects by considering the model on a finite-dimensional lattice such as a one-dimensional chain instead of the complete graph.
+Based on finite-dimensional lattice models, a mixing process could also be introduced as a parameter to study how such lattice models go through a transition to the corresponding population dynamics such as a Lotka-Volterra type of equation FMO .
+The relationship between lattice models and population dynamics is one topic discussed in the context of intermediate disturbance hypothesis Svensson12 .
+Those cases remain to be studied.
+Acknowledgements
+The authors thank Obihiro University of Agriculture and Veterinary Medicine for providing the facilities to accomplish this work.
+Appendix A The equations to determine the stationary solutions of the population dynamics
 As mentioned in the main text, based on the model we study, we consider ${\bm{Q}}:=(Q_{1},Q_{2},Q_{3})\in[0,1]^{3}$ where $Q_{h}$ is the density of each state $h$.
@@
 $\frac{dQ_{h}}{dt}=V_{h}({\bm{Q}}),$ (33)
-$V_{h}({\bm{Q}})=-\delta_{h,3}rQ_{3}+Q_{2}U\Bigl{(}2\to h\mid{\bm{Q}}\Bigr{)}+\delta_{h,1}r\sum_{h^{\prime}=2}^{3}Q_{h^{\prime}}-Q_{h}U\Bigl{(}h\to 2\mid{\bm{Q}}\Bigr{)}.$
+$V_{h}({\bm{Q}})=-\delta_{h,3}rQ_{3}+Q_{2}U\Bigl{(}2\to h\mid{\bm{Q}}\Bigr{)}$
+$+\delta_{h,1}r\sum_{h^{\prime}=2}^{3}Q_{h^{\prime}}-Q_{h}U\Bigl{(}h\to 2\mid{\bm{Q}}\Bigr{)}.$
 We define $Q_{h}^{*}$ such that $V_{h}({\bm{Q}}^{*})=0$ holds for $h=1,3$.
-Let us remind that because of $Q_{1}^{*}+Q_{2}^{*}+Q_{3}^{*}=1$, we have
+Let us remember that because of $Q_{1}^{*}+Q_{2}^{*}+Q_{3}^{*}=1$, we have
 $Q_{2}^{*}=1-Q_{1}^{*}-Q_{3}^{*},$ (34)
@@
 $G_{\beta}^{(h)}(a)=\tanh\beta\Big{(}(\delta_{h,1}-\delta_{h,3})2a+1\Big{)}.$ (37)
-I.1 The explicit expressions of $Q_{h}^{*}$ and $\chi_{h}$ at $r\to 0$
+A.1 The explicit expressions of $Q_{h}^{*}$ and $\chi_{h}$ at $r\to 0$
 Based on (36) for $T>0$, we find the stationary solution of the densities $Q_{1}^{*}$ and $Q_{3}^{*}$ satisfying the following:
@@
 $Q_{1}^{*}(T,0)=Q_{3}^{*}(T,0)=\frac{1}{2+\exp(-2\beta)},$ (39)
-$Q_{2}^{*}(T,0)=\frac{\exp(-2\beta)}{2+\exp(-2\beta)}.$
+$Q_{2}^{*}(T,0)=\frac{\exp(-2\beta)}{2+\exp(-2\beta)}.$ (40)
 Next, we focus on the case of $T=0$ where the following holds:
 $\lim_{T\to 0}G_{\beta}^{(h)}(a)=\begin{cases}1&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}>0\\ -1&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}<0\\ 0&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}=0.\end{cases}$ (41)
-There are $6$ cases in terms of the possibilities of taking values of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$, each of which can be $-1$, 00, or $1$.
+There are six cases in terms of the possibilities of taking values of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$, each of which can be $-1$, 00, or $1$.
 Note that if one of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$ is 00, the value of the other is uniquely determined.
 Indeed, when the two conditions of
-$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1>0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (9)
+$-2Q_{1}^{*}+2Q_{3}^{*}+1>0,$ (42)
+$2Q_{1}^{*}-2Q_{3}^{*}+1>0,$ (43)
 hold, based on (36), we obtain
@@
 $\lim_{T\to 0}G_{\beta}^{(h)}(Q_{3}^{*}-Q_{1}^{*})=1,$ (45)
-for $h=1,3$. Additionally, the two conditions of (9) lead to
+for $h=1,3$. Additionally, the two conditions of (42) and (43) lead to
 $\frac{1}{4}<Q_{1}^{*}(0,0)<\frac{3}{4},$ (46)
-$Q_{3}^{*}(0,0)=1-Q_{1}^{*}(0,0),$
-$Q_{2}^{*}(0,0)=0.$
-Note that the other $5$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
+$Q_{3}^{*}(0,0)=1-Q_{1}^{*}(0,0),$ (47)
+$Q_{2}^{*}(0,0)=0.$ (48)
+Note that the other five cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of Eq. (36).
 In order to compute $\chi_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial r}$ for $T>0$, let us perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $r=0$ for $h=1,3$ as follows:
 $Q_{h}^{*}(T,r)=Q_{h}^{*}(T,0)+\chi_{h}(T,0)r+O(r^{2}).$ (49)
-Then, substituting this form with (7) into (4), for $O(\beta r)\ll 1$ up to $O(r)$ order, we obtain
+Then, substituting this form with (39) and (40) into (36), for $O(\beta r)\ll 1$ up to $O(r)$ order, we obtain
 $-2\chi_{1}-g_{\beta}\chi_{3}+\frac{4g_{\beta}^{\prime}\beta}{2+g_{\beta}}(\chi_{3}-\chi_{1})+\frac{8}{2+g_{\beta}}=0,$ (50)
@@
 $\chi_{1}(T,0)=4\frac{(2+g_{\beta})(4+g_{\beta}^{2})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$ (53)
-$\chi_{3}(T,0)=4\frac{-4g_{\beta}(2+g_{\beta})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$
+$\chi_{3}(T,0)=4\frac{-4g_{\beta}(2+g_{\beta})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$ (54)
 where $g_{\beta}^{\prime}\beta\simeq\beta\exp(-2\beta)\to 0$ as $\beta\to\infty$ and $g_{\beta}^{\prime}\beta\simeq\beta\to 0$ as $\beta\to 0$. Note that $\chi_{2}=-\chi_{1}-\chi_{3}$ by definition.
-I.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
-Among $6$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, let us consider the situation where the following two conditions hold for $T\ll 1$:
-$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1=0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0.\end{cases}$ (18)
+A.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
+Among six possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, let us consider the situation where the following two conditions hold for $T\ll 1$:
+$-2Q_{1}^{*}+2Q_{3}^{*}+1=0,$ (55)
+$0\displaystyle 2Q_{1}^{*}-2Q_{3}^{*}+1>$ 0 . (56)
 In this case, we need to take the limit of $T\to 0$ carefully. First, we can perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $T=0$ as
@@
 $\lim_{T\to 0}G_{\beta}^{(3)}(Q_{3}^{*}-Q_{1}^{*})\to 1.$ (60)
-Combining (21) with the first condition of (18) gives
+Combining (59) with the condition of (55) gives
 $Z_{1}=$ $Q_{1}^{*}(0,r)=-\frac{1}{4}r+\frac{3}{4},$ (61)
-$Z_{3}=$ $Q_{3}^{*}(0,r)=-\frac{1}{4}r+\frac{1}{4},$
-$Q_{2}^{*}(0,r)=\frac{r}{2}.$
-Let us move onto (20) into which we substitute (23), leading to
+$Z_{3}=$ $Q_{3}^{*}(0,r)=-\frac{1}{4}r+\frac{1}{4},$ (62)
+$Q_{2}^{*}(0,r)=\frac{r}{2}.$ (63)
+Let us move onto (58) into which we substitute (61) and (62), leading to
 $\tanh 2(-\mu_{1}+\mu_{3})$ $=\frac{-r^{2}-10r+3}{(r+3)(1-r)}=:\theta(r).$ (64)
 Since $-1<\tanh 2(-\mu_{1}+\mu_{3})<1$ should hold, $r$ satisfies $0<r<r_{0}=-3+2\sqrt{3}$ due to $\lim_{r\to 0}\theta(r)=1$ and $\lim_{r\to r_{0}}\theta(r)=-1$.
-Thus, it turns out that (23) as the stationary solution is valid only for $0<r<r_{0}$.
-Note that $\lim_{r\to 0}Q_{1}^{*}(0,r)=\sup Q_{1}^{*}(0,0)=\frac{3}{4}$, $\lim_{r\to 0}Q_{3}^{*}(0,r)=\inf Q_{3}^{*}(0,0)=\frac{1}{4}$, and $\lim_{r\to 0}Q_{2}^{*}(0,r)=Q_{2}^{*}(0,0)=0$ hold.
-Next, among the other $5$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, when the following two conditions
-$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1<0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (25)
+Thus, it turns out that (61), (62) and (63) as the stationary solution is valid only for $0<r<r_{0}$.
+Note that $\lim_{r\to 0}Q_{1}^{*}(0,r)=\sup Q_{1}^{*}(0,0)=\dfrac{3}{4}$, $\lim_{r\to 0}Q_{3}^{*}(0,r)=\inf Q_{3}^{*}(0,0)=\dfrac{1}{4}$, and $\lim_{r\to 0}Q_{2}^{*}(0,r)=Q_{2}^{*}(0,0)=0$ hold.
+Next, among the other five possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, when the following two conditions
+$-2Q_{1}^{*}+2Q_{3}^{*}+1<0,$ (65)
+$2Q_{1}^{*}-2Q_{3}^{*}+1>0,$ (66)
 hold, based on (36) we obtain
 $-\frac{1}{2}(1-r)Q_{1}^{*}+r(1-Q_{1}^{*})=0,$ (67)
-$\frac{1}{2}(1-r)(1-Q_{1}^{*}-Q_{3}^{*})-rQ_{3}^{*}=0,$
+$\frac{1}{2}(1-r)(1-Q_{1}^{*}-Q_{3}^{*})-rQ_{3}^{*}=0,$ (68)
 where we have used
@@
 $Q_{1}^{*}(0,r)=\frac{2r}{1+r},$ (71)
-$Q_{3}^{*}(0,r)=\frac{(1-r)^{2}}{(1+r)^{2}},$
-$Q_{2}^{*}(0,r)=\frac{2r(1-r)}{(1+r)^{2}}.$
-Substituting (29) into the two conditions of (25) gives
-$\left\{\begin{array}[]{l}3r^{2}+10r-1>0,\\ r^{2}+6r-3>0,\end{array}\right.$
+$Q_{3}^{*}(0,r)=\frac{(1-r)^{2}}{(1+r)^{2}},$ (72)
+$Q_{2}^{*}(0,r)=\frac{2r(1-r)}{(1+r)^{2}}.$ (73)
+Substituting (71) and (72) into the two conditions of (65) and (66) gives
+$3r^{2}+10r-1>0,$ (74)
+$r^{2}+6r-3>0,$ (75)
 which lead to $r_{0}<r\leq 1$.
-Thus, the obtained solution (29) is valid only for $r_{0}<r\leq 1$.
+Thus, the obtained solution of (71), (72), and (73) is valid only for $r_{0}<r\leq 1$.
 Combining two cases of smaller $r$ and larger $r$ than $r_{0}$, the left-sided limit and the right-sided limit of $Q_{h}^{*}$ as $r\to r_{0}$ coincide.
-Note that the other $4$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
+Note that the other four cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of Eq.
+(36).
 In order to compute the response $\mu_{h}(0,r)$, we can perform Taylor series expansion as
@@
 $\mu_{1}(0,r)=-\frac{1+r}{4}\text{artanh }\theta(r),$ (79)
-$\mu_{3}(0,r)=\frac{1-r}{4}\text{artanh }\theta(r).$
+$\mu_{3}(0,r)=\frac{1-r}{4}\text{artanh }\theta(r).$ (80)
 Note that $\mu_{2}=-\mu_{1}-\mu_{3}$ holds by definition, leading to
@@
 $\mu_{1}(0,r)=\mu_{3}(0,r)=\mu_{2}(0,r)=0.$ (83)
-I.3 The explicit expressions of $Q_{h}^{*}$ and $\nu_{h}$ at $\beta\to 0$
+A.3 The explicit expressions of $Q_{h}^{*}$ and $\nu_{h}$ at $\beta\to 0$
 Based on (36), taking into account
@@
 $\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})-\frac{(1-r)}{4}Q_{1}^{*}+r(1-Q_{1}^{*})$ $=0,$ (85)
-$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})-\frac{(1-r)}{4}Q_{3}^{*}-rQ_{3}^{*}$ $=0.$
+$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})-\frac{(1-r)}{4}Q_{3}^{*}-rQ_{3}^{*}$ $=0.$ (86)
 Solving these equations, we obtain
 $Q_{1}^{*}(\infty,r)=\frac{5r^{2}+10r+1}{3r^{2}+10r+3},$ (87)
-$Q_{3}^{*}(\infty,r)=\frac{(1-r)^{2}}{3r^{2}+10r+3},$
-$Q_{2}^{*}(\infty,r)=\frac{-3r^{2}+2r+1}{3r^{2}+10r+3},$
-which are shown in Fig. 1(a).
-(a)Refer to caption
-(b)Refer to caption
-Figure 1: (a) $Q_{h}^{*}$ as a function of $r$ for $T=\infty$. (b) $D_{\alpha}$ as a function of $r$ for $T=\infty$.
+$Q_{3}^{*}(\infty,r)=\frac{(1-r)^{2}}{3r^{2}+10r+3},$ (88)
+$Q_{2}^{*}(\infty,r)=\frac{-3r^{2}+2r+1}{3r^{2}+10r+3}.$ (89)
 In order to compute the response $\nu_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial\beta}$, we perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $T=\infty$ as
@@
 $-\frac{1-r}{4}(\nu_{1}+\nu_{3})-\frac{1-r}{4}\nu_{1}-r\nu_{1}$ $=0,$ (91)
-$-\frac{1-r}{4}(\nu_{1}+\nu_{3})-\frac{1-r}{4}\nu_{3}-r\nu_{3}$ $=0.$
+$-\frac{1-r}{4}(\nu_{1}+\nu_{3})-\frac{1-r}{4}\nu_{3}-r\nu_{3}$ $=0.$ (92)
 Solving these equations, we obtain
@@
 leading to $\nu_{2}=0$ by definition.
-Note that by taking the derivative of the stationary solutions from (39) with respect to $r$, we also have
+Note that by taking the derivative of the stationary solutions from (87), (88), and (89) with respect to $r$, we also have
 $\chi_{1}(\infty,0)=\frac{20}{9},$ (94)
 $\chi_{3}(\infty,0)=-\frac{16}{9},$ (95)
-$\chi_{2}(\infty,0)=-\frac{-4}{9},$ (45)
-which are consistent with the expression of (17) in the limit of $\beta\to 0$.
-II Hill numbers $D_{\alpha}(r)$ as a function of $r$
-Let us define $M_{\alpha}:=\log D_{\alpha}$ and ${\bm{Q}}^{*}:=(Q_{1}^{*},Q_{2}^{*},Q_{3}^{*})\in[0,1]^{3}$.
-Then, the sign of $\partial_{r}M_{\alpha}$ is the same as that of $\partial_{r}D_{\alpha}$.
-Indeed, we get the following:
-$\partial_{r}M_{\alpha}=\alpha(1-\alpha)^{-1}(\sum_{h}(Q_{h}^{*})^{\alpha})^{-1}F_{\alpha}({\bm{Q}}^{*}),$ (46)
-$F_{\alpha}({\bm{Q}}^{*})=\sum_{h}(Q_{h}^{*})^{\alpha-1}\chi_{h}.$ (47)
-Thus, the nontrivial part determining the sign of $\partial_{r}D_{\alpha}$ is $F_{\alpha}$, which has key information to find out the local maximum point $r_{\ell}(T,\alpha)$ of $D_{\alpha}$.
-Note that $F_{\alpha}=0$ at $\alpha=1$ holds due to $\sum_{h}\chi_{h}=0$.
-II.1 The existence of a local maximum of $D_{\alpha}$ at $T\to 0$ for any $0<\alpha<\infty$
-First, let us consider the case of $r<r_{0}$ and $T\to 0$ where we use equations (23) by putting $r=r_{0}$.
-Then, as shown in Fig.
-2(a), we obtain that $F_{\alpha}<0$ for $\alpha>1$, $F_{\alpha}>0$ for $10.199\leq\alpha<$ 1, and $F_{\alpha}<0$ for $\alpha\leq 0.198$ as $r\to r_{0}$.
-It means that for $\alpha\geq 0.199$, $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is positive, whereas it is negative for $0<\alpha\leq 0.198$.
-Second, let us consider the case of $r>r_{0}$ where we use equations (29) by putting $r=r_{0}$.
-Then, as shown in Fig.
-2(b), it turns out that $F_{\alpha}>0$ for $\alpha>1$ and $F_{\alpha}<0$ for $\alpha<1$.
-It means that $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is negative for any $\alpha>0$.
-Therefore, combining two cases together, we reach the conclusion that there is a local maximum point $r_{\ell}(0,\alpha)>0$ at $T\to 0$ for any $\alpha>0$.
-The concrete values of $r_{\ell}(0,\alpha)$ are $r_{0}$ for $\alpha\geq 0.199$, and get smaller than $r_{0}$ between $\alpha=0.198$ and $0.199$.
-Then, $\lim_{\alpha\to 0}r_{\ell}(0,\alpha)$ is numerically estimated as $0.451\cdots$.
-(a)Refer to caption
-(b)Refer to caption
-Figure 2: (a) $\lim_{r\to r_{0}}F_{\alpha}$ as a function of $\alpha$ for $r<r_{0}$. (b) $\lim_{r\to r_{0}}F_{\alpha}$ a function of $\alpha$ for $r>r_{0}$
-II.2 The absence of local maxima of $D_{\alpha}$ at $T\to\infty$ for $0<\alpha<\infty$
-Suppose the following condition that for a fixed $T$, $\chi_{1}>0$, $\chi_{2}<0$, $\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold for $0<r\leq 1$.
-On that condition, we can derive that $(Q_{1}^{*})^{\alpha-1}\chi_{1}$ as a term in $F_{\alpha}$ can be larger than $|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ for $\alpha>1$, and smaller than it for $10<\alpha<$ 1.
-This leads to that for $0<r\leq 1$, $F_{\alpha}>0$ holds for $\alpha>1$, and $F_{\alpha}<0$ holds for $10<\alpha<$ 1, corresponding to that there is no local maximum of $D_{\alpha}$ as a function of $r$ in $0<r\leq 1$ for $\alpha>0$.
-The derivation of the above conclusion is as follows. First, it is straightforward to see
-$|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ $=(Q_{2}^{*})^{\alpha-1}\chi_{1}+(Q_{3}^{*})^{\alpha-1}|\chi_{3}|-(Q_{2}^{*})^{\alpha-1}|\chi_{3}|$
-$=(Q_{3}^{*})^{\alpha-1}\chi_{1}+(Q_{2}^{*})^{\alpha-1}|\chi_{2}|-(Q_{3}^{*})^{\alpha-1}|\chi_{2}|,$ (48)
-where we have used $\sum_{h}\chi_{h}=0$. Hence, we obtain
-$\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}\leq|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|\leq\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}.$ (49)
-It means that for $\alpha>1$, the following holds:
-$\chi_{1}(Q_{1}^{*})^{\alpha-1}>\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (50)
-which immediately leads to $F_{\alpha}>0$. On the other hand, for $10<\alpha<$ 1, the following holds:
-$\chi_{1}(Q_{1}^{*})^{\alpha-1}<\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (51)
-which immediately leads to $F_{\alpha}<0$. Thus, we reach the conclusion mentioned above.
-For $T=\infty$, (39) directly means that the conditions of $\chi_{1}>0,\chi_{2}<0,\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold.
-Thus, it leads to the absence of local maxima of $D_{\alpha}$ in $0<r\leq 1$ for any $\alpha>0$ as shown in Fig.
-1(b).
-Note that the conditions mentioned above can be broken, in particular, for sufficiently small $T$, for which the above results are not applied.
-Nevertheless, using a nullcline method to compute the stationary solutions of (4), we have numerically estimated that there is $\alpha_{0}(T)>0$ such that for $\alpha>\alpha_{0}(T)$, there are no local maxima of $D_{\alpha}$ for $T\geq 0.26$, whereas for $T\leq 0.25$, there is a local maximum of $D_{\alpha}$ for any $\alpha>0$.
-III Monte Carlo sampling
-III.1 Updating rules
+$\chi_{2}(\infty,0)=-\frac{4}{9},$ (96)
+which are consistent with the expression of (53) and (54) in the limit of $\beta\to 0$.
+Appendix B Monte Carlo sampling
+Figure 8: The density of height, $\sum_{i}\delta_{x_{i},h}/N$, calculated by the Monte Carlo (MC) simulation with $N=1000$ and the solutions $Q_{h}$ of the population dynamics (33). $T=1$, $r=0.2$.
+B.1 Updating rules
+Figure 9: The variance $\sigma_{hh}$ and correlation coefficients $C_{hh^{\prime}}$ averaged over ten samples with different trajectories.
+$T=10^{-3}$, $t_{0}=100$, $\tau=150$, $N=10^{2},10^{3},10^{4}$.
 Let us explain how to obtain sample trajectories by Monte Carlo simulations and compare them with the solutions of the derived population dynamics (33).
@@
     Next, repeat the steps from $1$ to $3$:
-  2. 1.
-    Choose randomly a name $\ell\ (1\leq\ell\leq N)$ of an agent with uniform distribution. That is, the probability of taking each name is $1/N$.
-  3. 2.
+  2. (1)
+    Randomly choose a name $\ell\ (1\leq\ell\leq N)$ of an agent with uniform distribution. That is, the probability of taking each name is $1/N$.
+  3. (2)
     With the probability $r$, the chosen agent $\ell$ is disturbed, making the transition from state $x_{\ell}$ to $1$.
-  4. 3.
-    If the chosen agent $\ell$ is not disturbed, choose either positive or negative direction with probability $1/2$.
+  4. (3)
+    If the chosen agent $\ell$ is not disturbed, choose either a positive or negative direction with probability $1/2$.
     1. (a)
-      If positive direction is chosen and $x_{\ell}\neq 3$, make the transition from state $x_{\ell}$ to $x_{\ell}+1$ with the probability
-      $\frac{1}{2}\Bigg{(}1+\tanh{\beta\Big{(}2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})}+1\Big{)}\Bigg{)}.$ (52)
+      If a positive direction is chosen and $x_{\ell}\neq 3$, make the transition from state $x_{\ell}$ to $x_{\ell}+1$ with the probability
+      $\frac{1}{2}\Bigg{(}1+\tanh{\beta\Big{(}2(x_{\ell}-\frac{1}{N-1}\sum_{k\neq\ell}x_{k})}+1\Big{)}\Bigg{)}.$ (97)
     2. (b)
-      If negative direction is chosen and $x_{\ell}\neq 1$, make the transition from state $x_{\ell}$ to $x_{\ell}-1$ with the probability
-      $\frac{1}{2}\Bigg{(}1+\tanh\beta\Big{(}-2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})+1\Big{)}\Bigg{)}.$ (53)
-As shown in Fig. 3, we have confirmed that Monte Carlo simulations have good agreement with the solution of the population dynamics (1).
-Figure 3: The Monte Carlo (MC) simulation with $N=1000$ and the solutions of the population dynamics (1). $T=1$, $r=0.2$.
-III.2 Finite-size fluctuations
+      If a negative direction is chosen and $x_{\ell}\neq 1$, make the transition from state $x_{\ell}$ to $x_{\ell}-1$ with the probability
+      $\frac{1}{2}\Bigg{(}1+\tanh\beta\Big{(}-2(x_{\ell}-\frac{1}{N-1}\sum_{k\neq\ell}x_{k})+1\Big{)}\Bigg{)}.$ (98)
+As shown in Fig.
+8, we have confirmed that Monte Carlo simulations have good agreement with the solution of the population dynamics (33).
+Note that for convenience in the simulations, we have replaced $\frac{1}{N-1}\sum_{k\neq\ell}x_{k}$ by $\frac{1}{N}\sum_{k}x_{k}$; the caused errors are the order of $1/N$, which is negligible, at least, near stationary states for sufficiently large $N$, except for the vicinity of $r=1,T=0$.
+B.2 Finite-size fluctuations
 Let us characterize the finite size fluctuations in the density of each state by variance $\sigma_{hh}$, covariance $\sigma_{hh^{\prime}}$, and correlation coefficient $C_{hh^{\prime}}$ for $h,h^{\prime}\in\{1,2,3\}$, which are defined as
@@
 Indeed, for $r\leq 10^{-3}$ and $T=10^{-3}$, the dynamics gets very slow and the typical trajectories do not cross the stationary solutions by $t=t_{0}=100$.
-Figure 4: The variance $\sigma_{hh}$ and correlation coefficients $C_{hh^{\prime}}$ averaged over $10$ samples with different trajectories.
-$T=10^{-3}$, $t_{0}=100$, $\tau=150$, $N=10^{2},10^{3},10^{4}$.
 As shown in Fig.
-4, we have performed Monte Carlo simulations with different systems size $N$.
-If we straightforwardly extrapolate the behaviors for larger system size, there seem to be singular behaviors in the three types of variance and the three types of correlation function in the large size limit near $r=r_{0}$ for $r<r_{0}$.
+9, we have performed Monte Carlo simulations with different systems sizes $N$.
+If we straightforwardly extrapolate the behaviors for larger system sizes, there seem to be singular behaviors in the three types of variance and the three types of correlation function in the large size limit near $r=r_{0}$ for $r<r_{0}$.
 In particular, it is notable to point out that three correlation coefficients in the large size limit seem to have perfect correlation, corresponding to $|C_{hh^{\prime}}|=1$, for only $r<r_{0}$.
-Thus, the results obtained by Monte Carlo simulations suggest that there is a universal curve presumably independent of $N$.
-Further, in the larger system sizes, it is more obvious that $C_{hh^{\prime}}$ for $r<r_{0}$ show U-shaped curves.
+Additionally, the results obtained by Monte Carlo simulations suggest that $C_{hh^{\prime}}$ for $r<r_{0}$ tends to show U-shaped curves as the system size $N$ gets larger.
 It is notable that the bottoms of the U-shaped curves are close to $r=-5+2\sqrt{7}\simeq 0.291\cdots$, which is the point where $\mu_{h}$ changes its sign.